Several years ago, Rolen Riders issued preferred stock with a stated annual dividend of 8% of its $100 par value. Preferred stock of this type currently yields 8%. Assume dividends are paid annually.

What is the estimated value of Rolen's preferred stock? Round your answer to the nearest cent.

Suppose interest rate levels have risen to the point where the preferred stock now yields 14%. What would be the new estimated value of Rolen's preferred stock? Round your answer to the nearest cent.
Answer #1

Annual dividend=100*8%=8

Hence value of preferred stock=Annual dividend/Current yield

1.value of preferred stock=8/0.08=$100

2.value of preferred stock=8/0.14=$57.14(Approx).